OpenStudy (anonymous):

Odell tried to solve the equation 6(x+7)=4x-3 . He recorded the following steps. 6(x+7)=4x-3 6x=4x-10 2x=-10 x=-5 When Odell checked his solution it did not work. What was his mistake? Explain. Show the correct solution.

OpenStudy (anonymous):

he didnt multiply 6(x+7) right. should have been 6x+42

OpenStudy (anonymous):

yeah and i need to show the correct solution. can you help me do that? i tried to do it and it didnt work out...

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

for some reason i got that x was 22.5

OpenStudy (anonymous):

6(x+7)=4x-3 6x+42=4x-3 2x=-45 x=-45/2

OpenStudy (anonymous):

yeah and that equals -45/2 which is -22.5

OpenStudy (anonymous):

how do i know when to leave it as a fraction?

OpenStudy (anonymous):

well how do they ask you to leave your answer as?

OpenStudy (anonymous):

i dont know :/

OpenStudy (anonymous):

leave it as both to be safe. -45/2 and -22.5

OpenStudy (anonymous):

and how do i check the solution?

OpenStudy (anonymous):

plug in -22.5 for x and see if both sides are the same

OpenStudy (anonymous):

i tried that already , didnt turn out.

OpenStudy (anonymous):

i got it! if you subsitude the fraction -45/2 for x, its going to turn out. thank you <3

OpenStudy (anonymous):

hmmm okay 6(-22.5+7)=4(-22.5)-3 6(15.5)=-90-3 93=93

OpenStudy (anonymous):

yeah. you can use the fraction or the decimal and they will come out as the same answer

OpenStudy (anonymous):

oh okay :D thanks